OpenStudy (anonymous):

What is the inverse of the function f(x)=1/3x-2 a f-1(x)=3x-6 B f-1(x)=3x+6 C f-1(x)=-3x+6 D f-1(x)=-3x-6

OpenStudy (anonymous):

Put y = f(x) y = 1/3x-2 Solve for x

OpenStudy (anonymous):

How do I do that

OpenStudy (anonymous):

Isolate x on one side of the equation

OpenStudy (anonymous):

Can u solve it for me please

OpenStudy (anonymous):

Hey to find the inverse of functions you got to do 4 steps. 1. Substitute f(x) with y. 2. Reverse the x and y variables. 3. Solve for y. 4. Replace y with f^-1x

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Can you do the first step?

OpenStudy (anonymous):

Y=1/3x-3

OpenStudy (anonymous):

Everything's good but you accidentally wrote the last number wrong :) it's 2 :). Now can you do the 2nd step? ;)

OpenStudy (anonymous):

So i add 2 to the other side

OpenStudy (anonymous):

That's a bit later, but don't forget the 2nd step says to Reverse the x and y variables. Since our function is now y=1/3x-2, we'd switch the x & y over, and it'd now be x=1/3y-2 :) Do you get it so far? :)

OpenStudy (anonymous):

and yeah on the 3rd step you'd then add 2 to both sides, & multiply by 3 to both sides as you're solving for y.

OpenStudy (anonymous):

Yeah . do now lol I forgot about that part tthat's whats been screwing my up

OpenStudy (anonymous):

If you forget, just use this problem to help you and eventually you'll be fine :) Did you get the answer? :)

OpenStudy (anonymous):

No I'm confused again likeep how can u divided 1/3

OpenStudy (anonymous):

*like

OpenStudy (anonymous):

oh okay :) well we got x+2=1/3y To eliminate the 1/3, if we multiply it by it's reciprocal (opposite) which is 3, the 1/3 & 3 will cancel out leaving the y by itself :) and then we'd multiply 3 to the other side as well :) did you get that? :)

OpenStudy (anonymous):

you multiply instead to cancel out the fraction by it's opposite number :)

OpenStudy (anonymous):

Ohh pellet . Feel stupid now ugh

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

dam

OpenStudy (anonymous):

its okay :) everyone makes mistakes so it's a good thing you know how to :) I myself just learned this 2 days ago :p I made mistakes at first but I got it afterward :)

OpenStudy (anonymous):

Okeyyy thanks sooo much

OpenStudy (anonymous):

no prob. :) hey by some chance do you take algebra 2 on flvs? o.O

OpenStudy (anonymous):

No im tmaking it i. School some of it is eeasy but some of it is just plan out hard why

OpenStudy (anonymous):

oh okay, no prob. :) I take Alg. 2 on flvs so was just wondering bc just learned this as well :) well best of luck but i g2g now so bye :)

OpenStudy (anonymous):

OK bye